# Linear-time classical approximate optimization of cubic-lattice classical spin glasses

### v1

@@ Abstract
Computing low-energy configurations (i.e., approximate optimization) of classical spin glasses is of relevance to both condensed matter and combinatorial optimization. Recent theoretical work opens the possibility to make its time complexity with quantum annealing generically polynomial, and D-Wave experiments can now achieve approximate optimization of cubic-lattice classical Ising spin glasses with spins. It is therefore timely to ask which short-range classical spin glasses are good candidates for demonstrating quantum advantage in the time complexity of approximate optimization. One intuition is to consider models with very rugged energy landscapes in configuration space, where optimization is extremely slow with state-of-the-art Monte-Carlo methods. However, here we provide evidence that short-range classical spin glasses may be approximately optimized in linear time and space with a very simple deterministic tensor-network heuristic regardless of ruggedness. On the cubic lattice with up to 505050 spins, we obtain energy errors of 3% for the model used in recent D-Wave experiments, and 5% for much harder planted-solution instances, all in less than one hour per instance. For cubic-lattice-Ising reductions of unweighted Max-Cut on random 3-regular graphs with up to 300 vertices, we find energy errors of 1% and approximation ratios of about 72-88%. These results inform the search for quantum advantage and suggest an efficient classical method for generating warm starts for other spin-glass optimization algorithms. Our algorithm is amenable to massive parallelization and may also allow for low-power, accelerated implementations with photonic matrix-multiplication hardware.
@@ I Introduction
Recent developments in quantum annealing make the simple-cubic-lattice model of Eq. (1) particularly relevant to the issue of quantum advantage: On the experimental front, D-Wave and collaborators have recently demonstrated the approximate optimization of such a model with over 5,000 spins with pure quantum annealing [5] and over 11,000 spins with hybrid quantum annealing [6]. On the theoretical side, the recent work in Ref. [7] presents the possibility of achieving generic polynomial time complexity for (heuristic) approximate optimization of classical Ising spin glasses with quantum annealing. Therefore, quantum advantage in the computational time complexity for approximate optimization of Eq. (1) may be possible in the near- to mid-term, but it is not clear for which instance classes of , if any; understanding where known classical algorithms are slow with this task can provide clues.
State-of-the-art Markov-Chain-Monte-Carlo algorithms are considered to be some of the best general classical heuristics for approximate and exact optimization of spin glass models. Consequently, (classical) computational hardness has become synonymous in the spin-glass literature with Monte-Carlo hardness (e.g., Ref. [8]). For this reason, Ref. [9] suggests that spin-glass models that display a steep increase in the autocorrelation time of parallel-tempering Monte Carlo as the temperature is lowered are good places to demonstrate quantum advantage. Also, Ref. [10] shows a time-complexity advantage of (error-corrected) quantum annealing over parallel tempering Monte Carlo (with isoenergetic cluster moves) for approximate optimization of a quasi-2D spin glass, and concludes this to be a demonstration of algorithmic quantum speedup. The hardness of certain spin glass instances for Monte-Carlo (and other) optimization algorithms that traverse the spin configuration space is understood to be due to the rugged energy landscape (i.e., a plethora of local minima) that arises at low temperatures in that space. Even machine-learning enhancements of Monte Carlo methods have been shown to be ineffective at overcoming the hardness from such rugged landscapes [11]. However, it is thought that quantum algorithms may be able to overcome this source of hardness via quantum tunneling [12, 13, 14].
@@ IV Summary and Discussion
The first problem class that we tested was the model used in recent D-Wave experiments. The quantum annealing D-Wave experiments in Ref. [5] were at a fixed size of over 5,000 spins, and they achieved a minimum energy error of about in a time of about 1 . The hybrid quantum-classical D-Wave experiments at a size of over 11,000 spins in Ref. [6] achieved a minimum energy error of about 0.1–1% in about one second of total quantum processing time. In comparison, our heuristic consistently yielded a minimum energy error of about at sizes 8,000–125,000 spins in a time – seconds. It may still be possible for quantum annealing to show a quantum advantage for approximate optimization in this problem class if it demonstrates either a sublinear time-complexity or a superlinear time-complexity with a minimum energy error of less than 3% consistently over a range of system sizes similar to the range in this work.
@@
The second problem class that we tested was the much more rugged instance class of the cubic-lattice tile planting model. Ref. [31] showed this problem class to have an exact-optimization time a few orders of magnitude larger than the model for the leading Monte Carlo heuristics at a system size of 888 spins. Here our heuristic consistently yielded a minimum energy error of about at 8,000 spins and about at 125,000 spins; the error monotonically decreased with system size, suggesting extra error at the system boundaries compared to the bulk. The TTS was the same as with the model: – seconds. These results cast doubt on the belief that all rugged-energy-landscape spin glasses are good candidates for demonstrating quantum advantage.
@@
The third problem class that we tested was the “two-level spin glass” of Barahona [4], which encodes the problem of unweighted Max-Cut on random 3-regular graphs. For this problem class our heuristic consistently yielded a minimum energy error of at system sizes ranging 7,500–270,000 spins, corresponding to random-graph sizes of between 50 and 300 vertices. The approximation ratio was about 0.72–0.88, and the TTS was – seconds. The time-complexity of our deterministic heuristic was verified to be quadratic in the number of vertices. In comparison, the well-known Goemans-Williamson approximation algorithm for Max-Cut [35], which often serves as a benchmark for other Max-Cut algorithms, has a guaranteed approximation ratio of and runs in polynomial time.
@@
While the solution quality of our heuristic on the random-3-regular-graph Max-Cut problem does not approach the typical solution quality of the Goemans-Williamson algorithm, it is not very poor either. This seems surprising given that our heuristic operates via a series of local approximate optimizations, whereas the problem is intrinsically nonlocal. We propose the following explanation: the bootstrapping nature of the algorithm allows the non-local correlations to propagate over the lattice as the spins are successively decimated. It would be interesting in future work to test how well our heuristic works for other such nonlocal-to-local reductions. We conjecture that solution quality would decrease with increasing density of the original graph.
@@
The computational cost of the present heuristic is dominated by matrix multiplications, for which specialized photonic hardware is under active development. Implementation of a heuristic of the present type with such hardware could yield substantially lower time and energy costs. However, it is unlikely that such hardware will be able to achieve beyond about 8 bits of numerical precision in the foreseeable future [42], whereas the results presented here were all with 64 bits of precision. Our (unshown) preliminary tests of the algorithm with float16 showed very poor results for cubic-lattice tile planting but good results for Barahona’s two-level spin glass. Whether or not the present heuristic can be modified to yield sufficiently good warm starts when limited to 8-bit precision is an important open question.
As mentioned, it is possible (and in our opinion very likely) that the heuristic presented in this work will not work well for short-range spin glasses that are reductions or embeddings of dense-graph combinatorial optimization problems. Besides, such optimization problems are more economically represented as dense-graph spin glasses. A tensor-network algorithm that can efficiently optimize dense-graph rugged-energy-landscape spin glasses is therefore desirable. Ref. [43] demonstrates that all-to-all coupled Ising spin glasses can sometimes be optimized by iterating over optimizations of randomly selected subsets of all-to-all coupled spins. In future work we plan to investigate such an approach with the tensor-network machinery that we have demonstrated here, such that the subsets of all-to-all coupled spins take the place of the local fragments in the present algorithm. Our (unshown) preliminary tests on small but very rugged instances of the all-to-all coupled Wishart planted ensemble [44] show good results for approximate optimization without decomposition into subsets. As with the heuristic that we demonstrated in the present work, such a tensor-network approach is straightforward to generalize to beyond two-body interactions and beyond binary variables.

### v2

@@ Abstract
Computing low-energy configurations (i.e., approximate optimization) of classical spin glasses is of relevance to both condensed matter and combinatorial optimization. Recent theoretical work opens the possibility to make its time complexity with quantum annealing generically polynomial, and D-Wave experiments can now achieve approximate optimization of cubic-lattice classical Ising spin glasses with spins. It is therefore timely to ask which short-range classical spin glasses are good candidates for demonstrating quantum advantage in the time complexity of approximate optimization. One intuition is to consider models with very rugged energy landscapes in configuration space, where optimization is extremely slow with state-of-the-art Monte-Carlo methods. However, here we provide evidence that short-range classical spin glasses may be approximately optimized in linear time and space with a very simple deterministic tensor-network heuristic regardless of ruggedness. On the cubic lattice with up to 505050 spins, we obtain energy errors of 3% for the model used in recent D-Wave experiments, and 5% for much harder, planted-solution instances, all in less than one hour per instance. For cubic-lattice-Ising reductions of unweighted Max-Cut on random 3-regular graphs with up to 300 vertices, we find energy errors of 1% and approximation ratios of about 72-88%. These results inform the search for quantum advantage and suggest an efficient classical method for generating warm starts for other spin-glass optimization algorithms. Our algorithm is amenable to massive parallelization and may also allow for low-power, accelerated implementations with photonic matrix-multiplication hardware.
@@ I Introduction
Recent developments in quantum annealing make the simple-cubic-lattice model of Eq. (1) particularly relevant to the issue of quantum advantage: On the experimental front, D-Wave and collaborators have recently demonstrated the approximate optimization of such a model with over 5,000 spins with pure quantum annealing [5] and over 11,000 spins with hybrid quantum annealing [6]. On the theoretical side, the recent work in Refs. [7, 8] presents the possibility of achieving generic polynomial time complexity for (heuristic) approximate optimization of classical Ising spin glasses with quantum annealing. Therefore, quantum advantage in the computational time complexity for approximate optimization of Eq. (1) may be possible in the near- to mid-term, but it is not clear for which instance classes of , if any; understanding where known classical algorithms are slow with this task can provide clues.
State-of-the-art Markov-Chain-Monte-Carlo algorithms are considered to be some of the best general classical heuristics for approximate and exact optimization of spin glass models. Consequently, (classical) computational hardness has become synonymous in the spin-glass literature with Monte-Carlo hardness (e.g., Ref. [9]). For this reason, Ref. [10] suggests that spin-glass models that display a steep increase in the autocorrelation time of parallel-tempering Monte Carlo as the temperature is lowered are good places to demonstrate quantum advantage. Also, Ref. [11] shows a time-complexity advantage of (error-corrected) quantum annealing over parallel tempering Monte Carlo (with isoenergetic cluster moves) for approximate optimization of a quasi-2D spin glass, and concludes this to be a demonstration of algorithmic quantum speedup. The hardness of certain spin glass instances for Monte-Carlo (and other) optimization algorithms that traverse the spin configuration space is understood to be due to the rugged energy landscape (i.e., a plethora of local minima) that arises at low temperatures in that space. Even machine-learning enhancements of Monte Carlo methods have been shown to be ineffective at overcoming this source of computational hardness [12]. However, it is thought that quantum algorithms may be able to overcome this source of hardness via quantum tunneling [13, 14, 15].
@@ IV Summary and Discussion
The first problem class that we tested was the model used in recent D-Wave experiments. The quantum annealing D-Wave experiments in Ref. [5] were at a fixed size of over 5,000 spins, and they achieved a minimum energy error of about in a time of about 1 . The hybrid quantum-classical D-Wave experiments at a size of over 11,000 spins in Ref. [6] achieved a minimum energy error of about 0.1–1% in about one second of total quantum processing time. In comparison, our classical heuristic consistently yielded a minimum energy error of about at sizes 8,000–125,000 spins in a time – seconds. Our results significantly constrain the space of possibilities for quantum annealing to show quantum advantage in the time complexity of approximate optimization in this problem class. To establish such a quantum advantage, it is now necessary (but not sufficient) for quantum annealing to either demonstrate a sublinear time-complexity, or a superlinear time-complexity with a minimum energy error of less than 3% consistently over a range of system sizes similar to the range in this work.
@@
The second problem class that we tested was one with an extremely rugged energy landscape: the instance class of the cubic-lattice tile planting model. Ref. [32] showed this problem class to have an exact-optimization time a few orders of magnitude larger than the model for the leading Monte Carlo heuristics at a system size of 888 spins. Here our heuristic consistently yielded a minimum energy error of about at 8,000 spins and about at 125,000 spins; the error monotonically decreased with system size, suggesting extra error at the system boundaries compared to the bulk. The TTS was the same as with the model: – seconds. These results cast doubt on the belief that rugged-energy-landscape spin glasses are good candidates for demonstrating quantum advantage, and constrain the space of possibilities for quantum annealing to demonstrate quantum advantage in this problem class similar to in the -model problem class.
@@
The third problem class that we tested was the “two-level spin glass” of Barahona [4], which is a local (cubic-lattice) spin glass that encodes the non-local problem of unweighted Max-Cut on random 3-regular graphs. For this problem class our heuristic consistently yielded a minimum energy error of at system sizes ranging 7,500–270,000 spins, corresponding to random-graph sizes of between 50 and 300 vertices. The approximation ratio was about 0.72–0.88, and the TTS was – seconds. The time-complexity of our deterministic heuristic was verified to be quadratic in the number of vertices. In comparison, the well-known Goemans-Williamson approximation algorithm for Max-Cut [36], which often serves as a benchmark for other Max-Cut algorithms, has a guaranteed approximation ratio of and runs in polynomial time.
@@
While the solution quality of our heuristic on the random-3-regular-graph Max-Cut problem does not approach the typical solution quality of the Goemans-Williamson algorithm, it is not very poor either. This seems surprising given that our heuristic operates via a series of local approximate optimizations, whereas the problem is intrinsically nonlocal. We propose the following explanation: the bootstrapping nature of the algorithm allows the non-local correlations to propagate over the lattice as the spins are successively decimated. It would be interesting in future work to test how well our heuristic works for other local spin-glass problems that encode non-local graph problems. We conjecture that solution quality would decrease with increasing density of the original graph.
@@
The computational cost of the present heuristic is dominated by matrix multiplications, for which specialized photonic hardware is under active development. Implementation of a heuristic of the present type with such hardware could yield substantially lower time and energy costs. However, it is unlikely that such hardware will be able to achieve beyond about 8 bits of numerical precision in the foreseeable future [43], whereas the results presented here were all with 64 bits of precision. Our (unshown) preliminary tests of our algorithm with float16 showed very poor results for cubic-lattice tile planting but good results for Barahona’s two-level spin glass. Whether or not the present heuristic can be modified to generally yield sufficiently good warm starts when limited to 8-bit precision is an important open question.
As mentioned, it is possible (and in our opinion very likely) that the heuristic presented in this work will not work well for short-range spin glasses that are reductions or embeddings of dense-graph combinatorial optimization problems. Besides, such optimization problems are more economically represented as dense-graph spin glasses. A tensor-network algorithm that can efficiently optimize dense-graph rugged-energy-landscape spin glasses is therefore desirable. Ref. [44] demonstrates that all-to-all coupled Ising spin glasses can sometimes be optimized by iterating over optimizations of randomly selected subsets of all-to-all coupled spins. In future work we plan to investigate such an approach with the tensor-network machinery that we have demonstrated here, such that the subsets of all-to-all coupled spins take the place of the local fragments in the present algorithm. Our (unshown) preliminary tests on small but very rugged instances of the all-to-all coupled Wishart planted ensemble [45] show good results for approximate optimization without decomposition into subsets. As with the heuristic that we demonstrated in the present work, such a tensor-network approach would be straightforward to generalize to beyond two-body interactions and beyond binary variables.
@@ Acknowledgements.
- Bernaschi et al. [2024] M. Bernaschi, I. González-Adalid Pemartín, V. Martín-Mayor, and G. Parisi, The quantum transition of the two-dimensional ising spin glass, Nature 631, 749 (2024).